OpenStudy (anonymous):

Perform the indicated operations: (x^2+x-6)/(x^2+4x-12)÷(x+3)/(x-1)

OpenStudy (campbell_st):

factor the 2 quadratics... and remember when you divide by a fraction, flip and multiply. then you should see common factors that can be removed.

OpenStudy (anonymous):

The answer that I got is: (x-1)/(x+6) after I factored and cancelled. What am I doing wrong? The answer is suppose to be: 1/(x+6)(x-1).

OpenStudy (campbell_st):

ok... so its \[\frac{(x +3)(x-2)}{(x +6)(x -2)} \times \frac{(x -1)}{(x + 3)}\] well I got the same as you...

OpenStudy (anonymous):

Maybe my teacher put the wrong answer then.

OpenStudy (anonymous):

Thank you for your help!